### v1

# A Thorough Assessment of the Non-IID Data Impact in Federated Learning

Abstract
Federated learning (FL) allows collaborative machine learning (ML) model training among decentralized clients’ information, ensuring data privacy. The decentralized nature of FL deals with non-independent and identically distributed (non-IID) data. This open problem has notable consequences, such as decreased model performance and larger convergence times. Despite its importance, experimental studies systematically addressing all types of data heterogeneity (a.k.a. non-IIDness) remain scarce. This paper aims to fill this gap by assessing and quantifying the non-IID effect through a thorough empirical analysis. We use the Hellinger Distance (HD) to measure differences in distribution among clients. Our study benchmarks five state-of-the-art strategies for handling non-IID data, including label, feature, quantity, and spatiotemporal skewness, under realistic and controlled conditions. This is the first comprehensive analysis of the spatiotemporal skew effect in FL. Our findings highlight the significant impact of label and spatiotemporal skew non-IID types on FL model performance, with notable performance drops occurring at specific HD thresholds. Additionally, the FL performance is heavily affected mainly when the non-IIDness is extreme. Thus, we provide recommendations for FL research to tackle data heterogeneity effectively. Our work represents the most extensive examination of non-IIDness in FL, offering a robust foundation for future research.
Index Terms:
Federated Learning, Machine Learning, non-IID data, data heterogeneity quantification.
publicationid: pubid: $0$–$00$ 2021 IEEE
I Introduction
In today’s digital age, the interaction of machine learning (ML) and healthcare or financial data holds immense promise for improving disease diagnosis [36] and combating financial crimes [35]. However, it raises critical questions about data privacy when dealing with sensitive information from hospitals or banks. In this context, trusted research environments emerge as a mechanism for balancing ML research and protecting individual privacy [13].
Federated learning (FL) [30] has emerged as a transformative approach for training ML models across decentralized data sources, preserving data privacy and security. However, a significant challenge inherent in FL is the variation in data distributions across clients referred to as non-IID (Non-Independent and Identically Distributed) data. This so-called non-IIDness hinders model performance and convergence during training [44].
Motivation.
Despite the growing body of research focusing on the challenges posed by non-IID data in FL, there remains a gap in experimental studies that systematically evaluate the effectiveness of FL algorithms under different types of data heterogeneity. Understanding the extent and nature of non-IID characteristics are crucial for developing robust, fair, and widely applicable FL algorithms. In this context, our study uses the Hellinger Distance (HD[12] to quantify the differences in client data distributions. This approach ensures that our conclusions are robust and generalizable across diverse scenarios.
Contribution.
The subsequent points encapsulate the contributions of our study:
  1. 1.
    We benchmark five of the most employed state-of-the-art aggregation and client selection of FL algorithms to tackle non-IID data distributions among clients under realistic, controlled, and quantifiable methods for synthetic partitioning and all non-IID types (label, feature, quantity, and spatiotemporal skewness). This is the first study thoroughly analyzing how the spatiotemporal skew affects the performance of FL models.
  2. 2.
    We motivate using HD to quantify the differences among data distributions, standardizing the guidelines for systematic studies of data heterogeneity in FL. This is the first work to demonstrate that the effect of the non-IIDness is not the same under all the levels of heterogeneity We use HD to quantify differences in distribution as it provides more granular information, and we leave as future work the exploration of other measures; see our section on design insights and opportunities.
  3. 3.
    We provide a reference to researchers about which methods are robust to which kind of non-IIDness on highly benchmarked datasets.
  4. 4.
    We give highlights and relevant recommendations for FL researchers based on quantifying non-IIDness.
To the best of our knowledge, this is the most comprehensive and complete study of non-IIDness and its effects on FL models.
II Related Work
Studies that analyze and benchmark the performance of methods to tackle the non-IIDness effect on the FL models under controlled and systematic scenarios are scarce. Nevertheless, in this section, we introduce those works that, to some extent, provide empirical analysis about the data heterogeneity repercussions.
A study by Vahidian et al. [43] challenges conventional thinking regarding data heterogeneity in FL. They posit that dissimilar data among participants is not always detrimental and can be advantageous, and we found similar results. Their argument centers on two main points: firstly, that differences in labels (label skew) are not the sole determinant of data heterogeneity, and secondly, that a more effective measure of heterogeneity is the angle between the data subspaces of participating clients. Conversely, we encompass a broader spectrum of non-IIDness types and include images and tabular data.
Wong et al. [45] conduct extensive experiments on a large network of IoT and edge devices to present FL real-world characteristics, including learning performance and operation (computation and communication) costs. Moreover, they mainly concentrate on heterogeneous scenarios, which is the most challenging issue of FL. Although they thoroughly analyze the impact of non-IIDness, they focus on image datasets and do not compare any aggregation method to solve the effects of highly heterogeneous scenarios. Contrarily, this work includes diverse datasets and benchmark state-of-the-art methodologies to handle non-IIDness in FL.
In their study, Mora et al. [31] examine existing solutions in the literature to mitigate the challenges posed by non-IID data. On the one hand, they emphasize the underlying rationale behind these alternative strategies and discuss their potential limitations. Ultimately, they identify the most promising approaches based on empirical results and critical defining characteristics, such as any assumptions made by each strategy. In parallel, they focus only on label skew and only consider one dataset in their experiments. In our paper, we alternatively analyze broader datasets and data skewness types, identifying broader limitations and potential approaches to overcome them.
Wong et al. [45] present a comprehensive analysis of FL characteristics in real-world environments, focusing on large networks of IoT and edge devices. Their work offers valuable insights into learning performance and operational costs, particularly in heterogeneous scenarios. While they provide a detailed exploration of non-IID data issues using image datasets, they do not extensively evaluate different aggregation methods to address highly heterogeneous scenarios. In contrast, our work broadens this scope by incorporating diverse datasets and benchmarking state-of-the-art methods specifically designed to tackle non-IID challenges in FL.
One of the novelties of our study is the use of metrics to quantify non-IIDness among clients. The latter facilitates the systematic and regulated selection of non-IID. To the best of our knowledge, our work is the first to employ these metrics in the relevant bibliography.
III Background
III-A Federated learning basics
FL [30] is particularly suited for scenarios with siloed data (clients, local nodes, parties) where multiple organizations or institutions hold their datasets. This decentralized approach enables collaborative model training without sharing the raw data, contributing to data privacy and ownership for each participating organization [10].
Figure 1: FL training process overview
Figure 1 presents an overview of the cross-silo FL training process, where participating clients train local models ($M_{i}$) on their datasets ($\mathcal{D}_{i}$), all based on a pre-distributed global model ($M$[38, 48]. Instead of sharing raw data, clients exchange model updates ($M_{i}$) without sensitive information, aggregating centrally to enhance the global model. By utilizing a central server for coordination, silos transmit their updates, aggregated to improve the global model. This iterative process allows collaboration without compromising data privacy, as each client receives the updated global model without sharing raw data, ensuring data privacy while facilitating collaboration.
Data Skewness Types.
A centralized dataset [^1] Notice that this definition of a centralized dataset includes tabular data, images, medical data, graph data, and any dataset expressible as a collection of arrays. $\mathcal{D}=\{(\mathbf{x}_{1},y_{1}),\dots,(\mathbf{x}_{n},y_{n})\}$, is a collection of $n$ tuples where $\mathbf{x}_{i}=[(x_{i})_{1},\dots,(x_{i})_{m}]$ is the feature representation of the $i$th element (sample) in the dataset, and $y_{i}\in\{1,\dots,\ell\}$ is the (true) label of the $i$th element.
In the FL setting, the dataset $\mathcal{D}$ is distributed over $K$ clients. We let $\mathcal{D}_{i}$ be the set of elements of the $i$th client. That is:
$\mathcal{D}=\cup_{i=1}^{K}\mathcal{D}_{i}\qquad\text{and for $i\neq j$:}\qquad\mathcal{D}_{i}\cap\mathcal{D}_{j}=\emptyset.$
Defining the type of non-IIDness in FL is relevant since it can drastically influence the performance of the models. We follow the settings of previous work [28, 50]. For a supervised learning task on client $i$ (local node $i$), we assume that each data sample $(x,y)\in\mathcal{D}_{i}$, where $x$ is the input attributes or features, and $y$ is the label, following a local distribution $P_{i}(\mathbf{x},y)$. Let us define:
$P_{i}^{Y}(y)=\sum_{\begin{subarray}{c}(\mathbf{x},z)\in\mathcal{D}_{i}\\ z=y\end{subarray}}P_{i}(\mathbf{x},z)\hskip 8.53581pt$ $P_{i}^{X_{\ell}}(x)=\sum_{\begin{subarray}{c}(\mathbf{x},y)\in\mathcal{D}_{i}\\ x_{\ell}=x\end{subarray}}P_{i}(\mathbf{x},y)$
with $P_{i}^{Y}(y)$, the $i$th client labels’ distribution and $P_{i}^{X_{\ell}}(x)$ the distribution over the $\ell$th input feature of the $i$th client. Then, the classification for non-IID data (i.e., data skewness types) is as follows:
  • Regarding the concept of identically distributed:
    1. 1.
      Label skew: Means that the label distribution $P_{i}^{Y}(y)$ of different clients is different.
    2. 2.
      Feature skew: Occurs when the distribution of the features $P_{i}^{X_{\ell}}(x)$ varies from client to client.
    3. 3.
      Quantity skew: Refers to the significant difference in the number of examples of different client data $P_{i}(\mathbf{x},y)$.
  • Regarding the concept of independent:
    1. 4.
      Spatiotemporal skewness: Refers to the inner correlation of data in the time (or space) domain. In other words, the distribution $P_{i}(\mathbf{x},y)$ is not stationary but depends on the time or space.
Quantifying the degree of non-IIDness.
Regarding the selection of valuable scenarios to demonstrate the effects of non-IIDness in FL, the current literature often relies on ad-hoc partitions [26, 21, 15]. Therefore, in this work, we use a metric that systematically evaluates the level of non-IID data to select scenarios for measuring the effect of non-IIDness. We opted for the Hellinger Distance (HD), a metric widely used to gauge the separation between two probability distributions calculated as in Equation 1 [12].
$\textsf{HD}(P_{1}^{Y}(y),P_{2}^{Y}(y))=\frac{1}{\sqrt{2}}\sqrt{\sum_{y\in Y}\left(\sqrt{P_{1}^{Y}(y)}-\sqrt{P_{2}^{Y}(y)}\right)^{2}}$
Aggregation and client selection algorithms
In an FL process, the server aggregates the weights obtained from each client and communicate them back to each participant. In this section, we explain the five state-of-the-art aggregation and client-selection algorithms assessed in our experiments.
FedAvg: It is a fundamental algorithm in FL [30] designed to train ML models across a network of decentralized devices while preserving data privacy. In FedAvg, each client computes model updates and sends them to a central server using local data. The server averages these updates to calculate a global model update and then sends it back to the clients. The latter process iterates until it convergences (the model’s performance gets stable).
FedProx: It is a framework designed to address heterogeneity in FL [24], offering a generalized and reparametrized version of FedAvg. This approach incorporates a regularization term ($\mu$) to minimize the difference between local and global weights. Finally, the framework aggregates local model updates from all devices to obtain an updated global model. Using this proximal term, FedProx aims to improve convergence and performance in heterogeneous federated learning environments.
Random size-proportional selection (Rand): Rand is a baseline client selection strategy introduced to handle non-IID data that does not bias toward clients with higher local losses. Most current analysis frameworks consider a scheme that selects the training set of clients $S(t)$ by sampling $m$ clients randomly (with replacement) such that client $k$ gets selected with probability $p_{k}$, the fraction of data at that client [6].
Power-Of-Choice (POC): The POC algorithm performs well under non-IID distribution. It is inspired by the power of $d$ choices load balancing strategy, which queueing systems commonly use [6][7]. The central server first samples a candidate set of $d$ clients, where $d$ is between $m$ (the number of clients to be selected) and $K$. These candidates are chosen based on their data fraction ($p_{k}$). The server then sends the current global model to these candidates, who compute and return their local losses. Finally, the server selects $m$ clients with the highest losses to participate in the next training round. This approach aims to balance the workload and prioritize clients with more informative updates, improving the efficiency of the FL process.
Model Contrastive Learning (MOON): MOON is a simple and effective FL framework designed to tackle non-IIDness. It uses the similarity between model representations to correct the local training of individual parties (i.e., conducting contrastive learning at the model level) [23]. The network proposed in MOON has three components: a base encoder, a projection head, and an output layer. The base encoder extracts representation vectors from inputs. Le et al. introduce an additional projection head to map the representation to a space with a fixed dimension. Last, the output layer produces predicted values for each class. For ease of presentation, with model weight $w$, they use $F_{w}(\cdot)$ to denote the whole network and $R_{w}(\cdot)$ to denote the network before the output layer (i.e., $R_{w}(X_{\ell})$ is the mapped representation vector of input $X_{\ell}$).
IV Experimentation Setup
Datasets
This work considers six widely employed datasets to train the CL and FL models. Four of them, i.e., CIFAR10 [19], FMNIST [46], Physionet 2020 [14], Covtype [4], serve to simulate label, feature, and quantity skew, while the remaining two datasets, i.e., 5G Network Traffic flows [8], Snapshot Serengeti [41] work for spatiotemporal skew simulation. Table I summarizes the main characteristics of the datasets.
Table I: Characteristics of the datasets
Dataset Type #training examples #test examples #features # classes Class distr.
CIFAR10 Images 50,000 10,000 3,072 10 Balanced
FMNIST Images 60,000 10,000 784 10 Balanced
Physionet Tabular 39,895 2,095 120 27 Balanced
Covtype Tabular 522,910 58,102 54 7 Unbalanced
Serengeti Tabular 257,927 28,659 64 13 Unbalanced
5G NTF Tabular 74,838 13,207 7 12 Unbalanced
Models
We adopt a well-studied convolutional neural network (CNN) broadly applied in computer vision [5] for the image datasets. It includes one input layer and three convolutional blocks—where the first two blocks each have a convolutional layer followed by a max pooling layer, and the final block contains a convolutional layer and a flattened layer. The initial convolutional layer has 32 filters, whereas the subsequent two layers each have 64 filters, all with a $3\times 3$ filter size and ReLu as the activation function. In the dense section of the network, there is one dense layer with 64 neurons using ReLu as the activation function. Additionally, we employ in our tests the transfer learning models EfficientNetB0 [42] and MobileNetV2 [39] since they produce higher classification power results for the datasets studied.
For the tabular datasets, we use a deep neural network (DNN), selected because it is widely employed in classification tasks with tabular data [37]. It comprises one input layer, three hidden layers, and one output layer [32]. The input layer uses as many units as the number of features used in the training set. The three layers contain 500 hidden units each, and the last layer is formed by considering the neurons equal to the number of classes to predict. Additionally, the hidden layers used the ReLu activation function and the output layer a SoftMax function. We use Adam as our optimizer with a learning rate of 0.001 for $K=30$ clients and a batch size of 64. We train the models for 40 communication rounds and ten local epochs for all aggregation algorithms except MOON, which is trained for 100 communication rounds. We run the experiments on all six datasets using five or ten different data partitions (i.e., five or ten random seeds).
IV-1 Hyperparameters tuning
For a fair comparison, we base our hyperparameter grids on the best-performing hyperparameters presented in the original papers as follows:
  • Rand: The fraction of clients considered in each communication round gets fine-tuned from $\{0.3,0.5,0.7\}$ [6].
  • FedProx: The $\mu$ parameter gets fine-tuned from {0, 0.001, 0.01, 0.1, 1, 10, 100} [24].
  • POC: The parameter $C$ is equal to 0.5. The parameter $d$ gets fine-tuned from $\{15,18,19,21\}$ [6].
  • MOON: The $\mu$ is tuned from the grid of $\{0.1,1,5,10\}$, and we find the best $\mu$ of 0.1, and we set the value of $temperature$ to 0.5 [23].
IV-A Methods for Synthetic Partitioning
The datasets introduced previously are all centralized, and we must split them while controlling the level of non-IIDness to train FL models. For such a purpose, we use the FedArtML [17] tool to partition each dataset among the clients. The following lines explain the partition method employed to simulate each skewness type.
IV-A1 Label Skew
We use the Dirichlet distribution (DD) to partition data among clients based on label distribution. The DD generates random numbers summing to one, controlled by parameter $\alpha$. Higher $\alpha$ values (e.g., 1000) create similar local distributions, while lower values increase the chance of clients having examples from a single, randomly chosen class [26]. Notice that the DD is the multivariate generalization of the Beta distribution, and the latter is the generalization of the Uniform distribution. Therefore, the partition of the datasets using DD is a skewed split of the data distribution [25].
Figure 2: Distribution of CIFAR10 among 30 clients for different levels of non-IIDness.
We quantify the level of non-IIDness of the data partitions on the clients using the HD. Our levels of non-IIDness, quantified using HD, would be approximately equal to the following values: $\{0.0,0.25,0.5,0.75,0.9\}$. Fig. 2 exemplifies the partition distribution for label skew using thirty clients. All ten classes get evenly distributed among every client in the IID scenario ($\alpha=1000,\textsf{HD}=0.0$). As we increase the $\alpha$ parameter in the DD, the distribution of classes among clients becomes more diverse. In the extreme case of $\alpha=0.03,\textsf{HD}=0.9$, certain classes are absent in some local nodes.
IV-A2 Feature Skew
For simulating feature skew, we employed two diverse methods from FedArtML to test their properties:
Gaussian noise method: This approach introduces diverse Gaussian noise levels to each client’s local dataset to achieve diverse feature distributions. Specifically, for each client $i$, noise levels $\hat{x}$ are added according to the user-defined noise level $\sigma$, with $\hat{x}\sim\text{Gau}\left(\sigma\cdot\frac{i}{K}\right)$, where $\hat{x}$ represents the resultant features after applying the noise level to the original features. Here, $\text{Gau}\left(\sigma\cdot\frac{i}{K}\right)$ denotes a Gaussian distribution with a mean of 0 and a variance of $\sigma\cdot\frac{i}{K}$, and $K$ represents the total number of clients.
Hist-Dirichlet-based method: The process starts by characterizing the attributes of each client using other average values and then subjecting them to a binning procedure. Subsequently, it establishes the participation of each feature category within each client using the DD with a specified $\alpha$. Unlike the Gaussian Noise approach, this method distributes the data among the clients without modifying the features.
We measure the non-IIDness with the HD among the features across clients (FHD) within the range $\{0,0.25,0.5,0.75,0.9\}$.
IV-A3 Quantity Skew
We employ the MinSize-Dirichlet method, which specifies the DD’s $\alpha$ value and generates the desired participation proportions for each client. Subsequently, a minimum required size, which we refer to as “the minimum number of examples,” is established for each client. Thus, the minimum proportion size, denoted as $MinSize$, is calculated as $\textit{MinSize}=\frac{\textit{MinRequiredSize}}{n}$, where $n$ represents the total number of examples in the centralized dataset. If the designated proportions fall below MinSize, it substitutes them with MinSize. Finally, the proportions are normalized to fall from 0 to 1.
We assess the level of non-IIDness using the HD for quantity skew (QHD) within the range $\{0,0.10,0.17\}$.
IV-A4 Spatiotemporal (SPT) skew
The primary constraint in this partition process is that the dataset must contain a categorical variable of space (i.e., places, cities, latitude, longitude, etc.) or time (i.e., time, months, years, etc.) to use as the partitioning variable. For instance, Fig. 3 depicts the distribution of labels along the date of the 5GNTF dataset. In this case, the space variable employed to create the federated data is the flow’s date expressed in year-month-day format (categorical).
Figure 3: Distribution of 5GNTF applications (label) along date (spatiotemporal variable expressed in YYYY-MM-DD).
We use the St-Dirichlet method, which employs the DD to segment the data based on spatial (SP skew) or temporal (TMP skew) categories to distribute the data among federated clients. We assess the level of non-IIDness using the HD for spatiotemporal skew (STHD) within the range {0, 0.25, 0.5, 0.75, 0.9}.
IV-A5 Hardware Specification
We used an Ubuntu 22.04.4 LTS machine with 200 GB of disk, Intel(R) Xeon(R) Platinum 8259CL CPU @ 2.50GHz processor, 16 processors, 125 GB of RAM, and Python 3.10.12 to run the experiments. The FL models were trained using the Flower [2].
V Results and Highlights
This section depicts the main highlights and results of the experiments defined in Section IV. Each result gets grouped by label, feature, quantity, and spatiotemporal.
V-A Label Skew
This subsection examines how label skewness in the client data affects the models’ performance. Consider that the Covtype is an unbalanced dataset regarding the labels, and CIFAR10, FMNIST, and Physionet are balanced datasets. The accuracy of the models on each centralized dataset is our baseline for comparing the accuracy of the models generated in FL.
V-A1 Classification power
In this subsection, we focus on the findings from the simulations to compare different aggregation algorithms and datasets regarding their classification power (a.k.a accuracy).
Table II: Mean and standard deviation Accuracy for each dataset for CL, FedAvg, Rand, FedProx, Power-Of-Choice, and MOON, considering different levels of non-IIDness as measured by HD for $K=30$. Each model has undergone ten different trials.
Category Dataset HD CL FedAvg Rand FedProx POC MOON
Label distribution skew CIFAR10 0 70.50% $\pm$ 0.60% 66.12% $\pm$ 0.73% 66.16% $\pm$ 0.74% 66.35% $\pm$ 0.72% 66.26% $\pm$ 0.70% 64.45% $\pm$ 1.05%
0.25 65.91% $\pm$ 0.49% 65.49% $\pm$ 0.52% 65.86% $\pm$ 0.60% 65.61% $\pm$ 0.67% 63.40% $\pm$ 0.74%
0.5 63.41% $\pm$ 0.95% 62.93% $\pm$ 1.55% 63.56% $\pm$ 0.83% 62.86% $\pm$ 1.71% 60.95% $\pm$ 1.33%
0.75 58.85% $\pm$ 1.06% 56.80% $\pm$ 2.24% 58.80% $\pm$ 1.04% 55.84% $\pm$ 1.48% 55.27% $\pm$ 0.51%
0.9 43.22% $\pm$ 2.24% 40.95% $\pm$ 2.43% 44.33% $\pm$ 2.83% 39.04% $\pm$ 2.99% 38.84% $\pm$ 2.31%
FMNIST 0 90.90% $\pm$ 0.20% 90.68% $\pm$ 0.18% 90.63% $\pm$ 0.18% 90.69% $\pm$ 0.15% 90.62% $\pm$ 0.17% 88.70% $\pm$ 0.27%
0.25 90.44% $\pm$ 0.14% 90.52% $\pm$ 0.17% 90.51% $\pm$ 0.17% 90.51% $\pm$ 0.18% 88.17% $\pm$ 0.22%
0.5 89.92% $\pm$ 0.11% 89.84% $\pm$ 0.31% 89.96% $\pm$ 0.20% 89.74% $\pm$ 0.35% 87.37% $\pm$ 0.22%
0.75 88.15% $\pm$ 0.54% 87.51% $\pm$ 0.81% 88.17% $\pm$ 0.47% 87.23% $\pm$ 0.78% 84.70% $\pm$ 0.84%
0.9 80.37% $\pm$ 3.78% 79.08% $\pm$ 4.67% 81.10% $\pm$ 2.21% 77.83% $\pm$ 3.28% 70.79% $\pm$ 5.73%
Physionet 0 63.74% $\pm$ 1.24% 57.97% $\pm$ 0.49% 57.48% $\pm$ 0.40% 58.16% $\pm$ 0.62% 57.86% $\pm$ 0.76% 61.80% $\pm$ 0.62%
0.25 57.65% $\pm$ 0.47% 57.42% $\pm$ 0.54% 57.79% $\pm$ 0.55% 57.48% $\pm$ 0.53% 60.94% $\pm$ 0.60%
0.5 55.69% $\pm$ 0.93% 55.26% $\pm$ 1.05% 56.29% $\pm$ 1.00% 55.24% $\pm$ 1.48% 58.76% $\pm$ 0.71%
0.75 50.88% $\pm$ 1.18% 50.19% $\pm$ 2.07% 51.47% $\pm$ 1.20% 49.51% $\pm$ 2.30% 53.51% $\pm$ 1.37%
0.9 41.35% $\pm$ 2.70% 39.68% $\pm$ 3.01% 41.95% $\pm$ 2.07% 38.81% $\pm$ 3.68% 42.49% $\pm$ 3.00%
Covtype 0 95.60% $\pm$ 0.10% 94.95% $\pm$ 0.06% 94.89% $\pm$ 0.08% 94.96% $\pm$ 0.09% 94.84% $\pm$ 0.10% 95.64% $\pm$ 0.06%
0.25 92.05% $\pm$ 1.14% 91.63% $\pm$ 1.52% 92.10% $\pm$ 1.28% 93.89% $\pm$ 0.73% 93.36% $\pm$ 1.18%
0.5 84.92% $\pm$ 3.71% 83.10% $\pm$ 2.79% 85.54% $\pm$ 3.39% 88.21% $\pm$ 2.17% 84.61% $\pm$ 3.44%
0.75 77.55% $\pm$ 3.63% 76.25% $\pm$ 4.10% 77.55% $\pm$ 3.64% 74.68% $\pm$ 5.31% 57.79% $\pm$ 12.66%
0.9 59.10% $\pm$ 8.70% 59.02% $\pm$ 9.19% 59.46% $\pm$ 8.74% 57.29% $\pm$ 10.72% 50.51% $\pm$ 5.57%
Number of times that performed the best 3 1 8 2 6
Highlight 1: The drop in the model’s performance appears in a double threshold. When HD is bigger than 0.5 and 0.75, a notable performance decline becomes immediately evident.
Previous works claim that the data non-IIDness affects the performance of FL models [27, 29, 16]. Nevertheless, for the first time, we demonstrate that the effect of non-IIDness is not the same under all the levels of heterogeneity. Fig. 4 depicts the accuracy change by varying the level of non-IIDness of the data distributions among the clients measured by HD concerning the baseline model created in the centralized setting. As the non-IIDness of the data partitions increases, the model’s accuracy decreases. When the HD between data distributions of the clients exceeds 0.75, the drop becomes more drastic compared to previous levels.
Figure 4: Changes in the models’ accuracy considering different levels of non-IIDness measured by HD for $K=30$.
Highlight 2: For high levels of non-IIDness the performance of transfer learning models gets more degraded
Figure 5: Changes in the models’ accuracy considering different levels of non-IIDness measured by HD for $K=30$ using CNN, EfficientNetB0 and MobilbeNetV2 on CIFAR10 dataset.
Fig. 5 illustrates the accuracy of models created using three different architectures: the CNN discussed earlier, EfficientNetB0, and MobileNetV2, both of which utilize transfer learning. The figure clearly shows a decline in model performance at the double thresholds after 0.5 and 0.75 HD. Notably, in pathological non-IID scenarios, the transfer learning models perform worse than our naive CNN model. This behavior can be attributed to the fact that only the last layers’ weights are updated locally, leading to skewed model weights and reducing the effectiveness of the aggregated model in capturing the overall dataset’s features, making transfer learning less suitable in such cases.
Highlight 3: Aggregation algorithms such as Rand, POC, and MOON are particularly vulnerable to performance degradation under conditions of high non-IIDness.
Consider the case when the HD is 0.9 (i.e., high non-IIDness) for all the datasets presented in Table II. When comparing the accuracy of Rand and POC versus their correspondent CL performance, those algorithms tend to have a higher drop in performance than the other aggregation methods. In FL scenarios with high non-IIDness, the distribution of data labels across clients is highly uneven. The latter means that specific clients may have more or different data types than others. Under such a scenario, Rand and POC may inadvertently select clients with skewed or unrepresentative data distributions, leading to poor generalization performance when aggregating their updates. MOON has the sharpest decrease in performance when the paradigm switches from CL to FL. Penalizing the current model with past and global representations made by each client will not help the model attain improved performance.
Highlight 4: Unbalanced class datasets experience a sharper drop in performance when moving from IID to highly non-IID settings compared to balanced datasets.
Table III: The performance’s decrease range of the model for the four datasets, moving from the lowest (HD=0) to the highest (HD=0.9) levels of non-IIDness and considering FedAvg.
Dataset (HD=0) - (HD=0.9)
CIFAR10 22.9%
FMNIST 10.31%
Physionet 16.62%
Covtype 35.85%
Consider the IID case (i.e., HD = 0) and the most extreme non-IID case (i.e., HD = 0.9) for each dataset as depicted in Table III. Notice that the range of decrease in the reported performance (accuracy of HD=0 - accuracy of HD=0.9) is higher for the unbalanced dataset (Covtype) than for balanced datasets (CIFAR10, FMNIST, and Physionet). The more significant decrease in performance for unbalanced datasets can relate to the challenge of learning from skewed and less representative data. In non-IID scenarios, where clients have different class distributions, models trained on these unbalanced datasets may struggle to generalize well across clients.
V-A2 Convergence
In this subsection, we focus on the findings obtained using the CIFAR10 dataset to compare different aggregation algorithms regarding their learning process and the smoothness of training.
Highlight 5: The higher the non-IIDness level of labels, the more required rounds to achieve convergence.
Table IV: The number of rounds required for FedAvg, Rand, FedProx, POC, and MOON to achieve 90% of the maximum accuracy reached in different levels of non-IID cases as determined by HD over CIFAR10 for K = 30.
Category Dataset Aggregation algorithm HD = 0 HD = 0.25 HD = 0.5 HD = 0.75 HD = 0.9
Label distribution skew CIFAR10 FedAvg 5 5 6 9 15
Rand 5 5 6 10 14
FedProx 5 5 6 9 15
POC 5 5 6 8 15
MOON 8 8 10 12 20
Table IV examines the algorithms’ convergence from a different perspective. We investigate the performance of each aggregation approach across a certain level of non-IID scenario independent from the others. For each specific non-IID situation described by HD, we determine how many rounds each aggregation algorithm requires to achieve 90% of its maximum accuracy. Therefore, regardless of the aggregation algorithm, as the non-IIDness of the data partitioning over the clients increases, more communication rounds are necessary to reach a convergence point (where the accuracy gets stable). We observe the latter because the data distributions among clients are sufficiently dissimilar, so the models built for each client are only optimal about their data, which diverges the model from the optimal case. As the training progresses, the weights delivered by the server to the clients improve because they get optimized by considering all of the data across all clients.
Table V: Mean and standard deviation Accuracy for each dataset for CL, FedAvg, Rand, FedProx, Power-Of-Choice, and MOON, considering different levels of feature skewness measured by FHD for K = 30. Each model has undergone ten different trials.
Category Method Dataset FHD CL FedAvg Rand FedProx POC MOON
Feature distribution skew Gaussian Noise CIFAR10 0 70.50% $\pm$ 0.60% 66.12% $\pm$ 0.70% 66.16% $\pm$ 0.74% 66.35% $\pm$ 0.72% 66.26% $\pm$ 0.70% 64.45% $\pm$ 1.05%
0.35 70.81% $\pm$ 0.45% 66.18% $\pm$ 0.57% 66.04% $\pm$ 0.44% 66.29% $\pm$ 0.63% 66.24% $\pm$ 0.50% 64.23% $\pm$ 0.53%
0.75 70.44% $\pm$ 0.29% 66.17% $\pm$ 0.44% 66.31% $\pm$ 0.61% 66.36% $\pm$ 0.56% 66.27% $\pm$ 0.56% 64.58% $\pm$ 0.55%
0.9 69.71% $\pm$ 0.03% 65.81% $\pm$ 0.82% 65.89% $\pm$ 0.72% 65.85% $\pm$ 0.68% 65.90% $\pm$ 0.77% 63.52% $\pm$ 0.66%
FMNIST 0 90.90% $\pm$ 0.20% 90.68% $\pm$ 0.18% 90.57% $\pm$ 0.14% 90.69% $\pm$ 0.15% 90.62% $\pm$ 0.17% 88.70% $\pm$ 0.27%
0.35 90.91% $\pm$ 0.26% 90.69% $\pm$ 0.19% 90.97% $\pm$ 0.15% 90.71% $\pm$ 0.17% 90.97% $\pm$ 0.17% 88.67% $\pm$ 0.28%
0.75 90.96% $\pm$ 0.11% 90.54% $\pm$ 0.09% 90.53% $\pm$ 0.11% 90.57% $\pm$ 0.23% 90.51% $\pm$ 0.19% 88.64% $\pm$ 0.19%
0.9 90.76% $\pm$ 0.13% 90.59% $\pm$ 0.09% 90.71% $\pm$ 0.15% 90.70% $\pm$ 0.29% 90.51% $\pm$ 0.11% 88.55% $\pm$ 0.21%
Physionet 0 63.74% $\pm$ 1.24% 57.97% $\pm$ 0.49% 57.48% $\pm$ 0.40% 58.16% $\pm$ 0.62% 57.86% $\pm$ 0.76% 61.80% $\pm$ 0.62%
0.35 63.30% $\pm$ 1.31% 57.89% $\pm$ 0.39% 57.92% $\pm$ 0.64% 58.08% $\pm$ 0.61% 57.47% $\pm$ 1.13% 61.22% $\pm$ 0.72%
0.75 60.13% $\pm$ 1.11% 52.81% $\pm$ 0.83% 52.84% $\pm$ 0.74% 52.28% $\pm$ 0.38% 51.39% $\pm$ 0.25% 56.10% $\pm$ 0.90%
0.9 28.97% $\pm$ 3.22% 29.43% $\pm$ 1.49% 29.88% $\pm$ 1.27% 29.43% $\pm$ 1.15% 25.25% $\pm$ 1.97% 32.06% $\pm$ 1.30%
Covtype 0 95.60% $\pm$ 0.10% 94.95% $\pm$ 0.06% 94.89% $\pm$ 0.08% 94.96% $\pm$ 0.09% 94.84% $\pm$ 0.10% 95.53% $\pm$ 0.04%
0.35 95.68% $\pm$ 0.10% 94.94% $\pm$ 0.06% 94.90% $\pm$ 0.04% 94.90% $\pm$ 0.08% 94.88% $\pm$ 0.08% 95.65% $\pm$ 0.06%
0.75 95.53% $\pm$ 0.04% 94.79% $\pm$ 0.08% 94.62% $\pm$ 0.04% 94.74% $\pm$ 0.07% 94.68% $\pm$ 0.13% 95.11% $\pm$ 0.07%
0.9 68.53% $\pm$ 1.49% 50.01% $\pm$ 0.35% 49.81% $\pm$ 0.50% 50.03% $\pm$ 0.42% 50.10% $\pm$ 1.67% 49.20% $\pm$ 0.11%
Hist-Dirichlet CIFAR10 0 70.50% $\pm$ 0.60% 66.42% $\pm$ 0.34% 66.35% $\pm$ 0.35% 66.21% $\pm$ 0.59% 66.40% $\pm$ 0.70% 64.79% $\pm$ 0.32%
0.25 66.09% $\pm$ 0.49% 66.13% $\pm$ 0.48% 65.82% $\pm$ 0.46% 66.15% $\pm$ 0.53% 65.12% $\pm$ 0.85%
0.5 66.30% $\pm$ 0.67% 66.04% $\pm$ 0.66% 66.22% $\pm$ 0.40% 66.52% $\pm$ 0.56% 64.52% $\pm$ 1.08%
0.75 66.23% $\pm$ 0.33% 66.04% $\pm$ 0.67% 66.17% $\pm$ 0.55% 66.34% $\pm$ 0.51% 64.99% $\pm$ 0.31%
0.9 66.25% $\pm$ 0.47% 65.25% $\pm$ 0.51% 65.25% $\pm$ 0.90% 66.17% $\pm$ 0.47% 64.12% $\pm$ 0.56%
FMNIST 0 90.90% $\pm$ 0.20% 90.72% $\pm$ 0.20% 90.68% $\pm$ 0.12% 90.68% $\pm$ 0.15% 90.59% $\pm$ 0.22% 88.75% $\pm$ 0.13%
0.25 90.62% $\pm$ 0.10% 90.66% $\pm$ 0.27% 90.70% $\pm$ 0.15% 90.62% $\pm$ 0.18% 88.72% $\pm$ 0.32%
0.5 90.78% $\pm$ 0.12% 90.66% $\pm$ 0.09% 90.63% $\pm$ 0.17% 90.78% $\pm$ 0.23% 88.43% $\pm$ 0.30%
0.75 90.53% $\pm$ 0.24% 90.67% $\pm$ 0.18% 90.73% $\pm$ 0.17% 90.56% $\pm$ 0.12% 88.26% $\pm$ 0.30%
0.9 89.77% $\pm$ 0.32% 89.73% $\pm$ 0.35% 89.66% $\pm$ 0.24% 89.81% $\pm$ 0.29% 87.38% $\pm$ 0.16%
Physionet 0 63.74% $\pm$ 1.24% 57.73% $\pm$ 0.72% 57.76% $\pm$ 0.51% 58.02% $\pm$ 0.76% 57.52% $\pm$ 0.68% 61.13% $\pm$ 0.42%
0.25 57.67% $\pm$ 0.61% 57.62% $\pm$ 0.69% 58.16% $\pm$ 0.63% 57.05% $\pm$ 0.37% 61.91% $\pm$ 0.39%
0.5 57.92% $\pm$ 0.40% 57.50% $\pm$ 0.68% 57.86% $\pm$ 0.33% 57.35% $\pm$ 0.47% 61.20% $\pm$ 0.83%
0.75 57.27% $\pm$ 0.64% 57.18% $\pm$ 0.97% 57.34% $\pm$ 0.88% 56.99% $\pm$ 0.86% 62.11% $\pm$ 0.44%
0.9 56.47% $\pm$ 0.80% 55.49% $\pm$ 1.34% 56.49% $\pm$ 0.60% 55.80% $\pm$ 1.24% 59.82% $\pm$ 1.02%
Covtype 0 95.60% $\pm$ 0.10% 94.95% $\pm$ 0.03% 94.81% $\pm$ 0.02% 95.00% $\pm$ 0.03% 98.84% $\pm$ 0.09% 95.62% $\pm$ 0.11%
0.25 94.95% $\pm$ 0.05% 94.83% $\pm$ 0.09% 94.97% $\pm$ 0.09% 94.77% $\pm$ 0.03% 95.63% $\pm$ 0.05%
0.5 94.90% $\pm$ 0.02% 94.88% $\pm$ 0.11% 94.90% $\pm$ 0.07% 94.88% $\pm$ 0.02% 95.65% $\pm$ 0.02%
0.75 94.80% $\pm$ 0.05% 94.67% $\pm$ 0.10% 94.78% $\pm$ 0.08% 94.74% $\pm$ 0.10% 95.57% $\pm$ 0.07%
0.9 93.30% $\pm$ 0.42% 93.11% $\pm$ 0.52% 93.38% $\pm$ 0.31% 93.47% $\pm$ 0.33% 94.51% $\pm$ 0.07%
Number of times that performed the best 4 2 7 7 16
Furthermore, FedAvg, Rand, FedProx, and POC exhibit similar behavior in achieving a convergence point, and they do so after roughly the same number of communication rounds. On the other hand, MOON, as expected, requires more rounds to reach the same convergence state.
V-B Feature Skew
This section examines how feature skewness in the client data affects the models’ performance.
V-B1 Classification power
In this subsection, we focus on the findings from the simulations to compare different aggregation techniques and datasets regarding their classification power (a.k.a accuracy) in the presence of feature skew over the clients’ data. Table V summarizes the models’ performance derived from different aggregation algorithms under varying degrees of non-IID feature distributions, as indicated by FHD.
Highlight 6: The performance of FL-generated models is lower than CL-generated models.
Transitioning the training methodology from CL to FL depicts a decline in the model’s performance, notably more pronounced in image datasets (CIFAR10) compared to tabular datasets (Covtype). This outcome is predictable since the model gets trained without access to the complete dataset, and each client optimizes the weights based on its available data. The more significant decrease observed in image datasets stems from the heightened complexity inherent in classification tasks compared to tabular datasets.
Highlight 7: The model’s performance in image datasets remains unaffected by increasing the level of feature non-IIDness.
Consider only the image datasets (CIFAR10, FMNIST) and the models generated in FL. Regardless of the aggregation algorithm employed, the performance of the final model remains stable across different levels of feature non-IIDness, consistently converging to specific values for each aggregation method. The latter pattern could occur due to convolutional layers in image analysis, where alterations in specific pixel values have minimal impact on the outcome of the convolution process in the Gaussian-noise method.
Highlight 8: For tabular datasets, using Gaussian noise levels that exceed FHD=0.9 results in a notable performance decline in the model, emphasizing the acute dissimilarity among samples.
Having tabular datasets (Covtype, Physionet) and using the Hist-Dirichlet approach shows that increasing the degree of feature non-IIDness does not impact the performance. However, when dealing with Gaussian noise, if we surpass FHD=0.9, there’s a noticeable decline in performance. This decline occurs because the data become so dissimilar (too noisy) from each other that even within the CL learning, the model struggles to learn from the data.
Highlight 9: In scenarios where the features are non-IID partitioned across clients, MOON demonstrates superior performance compared to all other aggregation algorithms for tabular datasets.
Table V validates that no particular algorithm outperforms others in image datasets, as they yield comparable final performance metrics. MOON emerges as the top-performing algorithm in both versions in tabular datasets, surpassing all other algorithms. Its performance is nearly equivalent to that of models trained in CL. The latter occurs since MOON can immediately start learning meaningful contrasts between differences in the label using the provided features of the tabular dataset. On the other hand, for images, the model first needs to learn to extract meaningful features from raw pixels before it can start contrasting different object classes effectively. For example, consider the Physionet dataset, which contains features such as age, sex, heart rate, and P-R interval. In that case, each feature has a clear medical interpretation, and the model can directly use the mentioned feature values without learning initial representations. Conversely, for CIFAR10, the model must learn to extract meaningful features from raw pixels before contrastive learning can be effective.
V-C Quantity Skew
This subsection examines how quantity skewness in the client data affects the models’ performance.
Table VI: Mean and standard deviation Accuracy for each dataset for CL, FedAvg, Rand, FedProx, Power-Of-Choice, and MOON, considering different levels of non-IID partitioning of the record quantities measured by QHD for $K=30$. Each model has undergone five different trials.
Category Method Dataset CL QHD FedAvg Rand FedProx POC MOON
Quantity distribution skew Min-size Dirichlet CIFAR10 70.50% $\pm$ 0.6% 0 65.77% $\pm$ 0.57% 65.92% $\pm$ 0.72% 66.45% $\pm$ 0.61% 66.04% $\pm$ 0.35% 64.01% $\pm$ 0.46%
0.10 66.69% $\pm$ 0.72% 66.29% $\pm$ 0.67% 66.71% $\pm$ 0.76% 68.82% $\pm$ 0.89% 63.24% $\pm$ 0.37%
0.17 66.04% $\pm$ 0.65% 67.91% $\pm$ 0.46% 68.07% $\pm$ 0.42% 68.44% $\pm$ 0.51% 63.49% $\pm$ 1.25%
FMNIST 90.90% $\pm$ 0.02% 0 90.72% $\pm$ 0.23% 90.69% $\pm$ 0.25% 90.64% $\pm$ 0.15% 90.67% $\pm$ 0.11% 88.49% $\pm$ 0.24%
0.10 90.37% $\pm$ 0.16% 90.39% $\pm$ 0.22% 90.30% $\pm$ 0.45% 90.78% $\pm$ 0.13% 88.04% $\pm$ 0.27%
0.17 90.39% $\pm$ 0.32% 90.43% $\pm$ 0.19% 90.44% $\pm$ 0.27% 90.65% $\pm$ 0.31% 88.10% $\pm$ 0.26%
Physionet 63.74% $\pm$ 1.24% 0 58.23% $\pm$ 0.68% 57.13% $\pm$ 0.56% 58.04% $\pm$ 0.29% 57.08% $\pm$ 0.53% 61.29% $\pm$ 0.81%
0.10 59.48% $\pm$ 2.09% 59.06% $\pm$ 1.71% 59.42% $\pm$ 1.23% 61.29% $\pm$ 0.50% 58.67% $\pm$ 1.97%
0.17 64.22% $\pm$ 1.27% 64.40% $\pm$ 0.55% 64.92% $\pm$ 0.71% 64.82% $\pm$ 0.85% 63.86% $\pm$ 0.40%
Covtype 95.60% $\pm$ 0.1% 0 94.97% $\pm$ 0.04% 94.87% $\pm$ 0.06% 94.96% $\pm$ 0.07% 94.83% $\pm$ 0.05% 95.15% $\pm$ 0.09%
0.10 95.67% $\pm$ 0.10% 95.65% $\pm$ 0.07% 95.70% $\pm$ 0.10% 95.79% $\pm$ 0.10% 95.13% $\pm$ 0.12%
0.17 90.65% $\pm$ 1.57% 94.77% $\pm$ 0.43% 95.27% $\pm$ 0.20% 94.94% $\pm$ 0.44% 94.23% $\pm$ 0.47%
Number of times that performed the best 1 0 3 6 2
V-C1 Convergence
In this subsection, we concentrate on the results of the simulations, aiming to contrast various aggregation methods and datasets in terms of their convergence.
Highlight 10: Feature skew doesn’t alter model convergence point
Table VII: The number of rounds required for FedAvg, Rand, FedProx, POC, and MOON to achieve 90% of the maximum accuracy reached in different levels of feature non-IID cases as determined by FHD over CIFAR10 and Covtype using Hist Dirichlet for K = 30
Category Method Dataset Aggregation algorithm FHD = 0 FHD = 0.25 FHD = 0.50 FHD = 0.75 FHD = 0.9
Feature distribution skew Hist Dirichlet CIFAR10 FedAvg 5 5 5 5 4
Rand 5 5 5 5 4
FedProx 5 5 5 5 4
POC 5 5 5 5 4
MOON 8 7 7 8 7
Covtype FedAvg 3 3 3 3 4
Rand 3 3 3 3 4
FedProx 3 3 3 3 4
POC 3 3 3 3 4
MOON 4 4 4 4 5
Table VII presents an alternative viewpoint on the latter concept. It outlines the iterations needed for FedAvg, Rand, FedProx, POC, and MOON to achieve 90% of the maximum accuracy across various degrees of feature non-IID conditions, as characterized by FHD, using the CIFAR10 dataset. Increasing the non-IIDness of features within the data has minimal impact on the model’s ability to converge to its optimal performance.
V-C2 Classification power:
In the following paragraphs, we concentrate on the simulation results to evaluate various aggregation methods and datasets regarding their classification accuracy, particularly considering the impact of quantity skew on the clients’ data.
Highlight 11: The presence of quantity skewness in the client’s data does not affect the performance of the final model.
Examining Table VI and considering each aggregation algorithm separately, it is evident that the performance of the final models remains consistent across various levels of quantity skewness in the clients’ records. This phenomenon occurs regardless of the chosen aggregation algorithm, confirming that quantity skewness does not affect model performance.
V-C3 Convergence
In this subsection, we focus on the findings obtained when different aggregation algorithms regarding their learning process and the smoothness of training.
Highlight 12: All aggregation algorithms converge after the same number of communication rounds in the presence of quantity skew.
Table VIII: The number of rounds required for FedAvg, Rand, FedProx, POC, and MOON to achieve 90% of the maximum accuracy reached in different levels of quantity non-IID cases as determined by QHD over CIFAR10 and Covtype using Min-size Dirichlet method for K = 30
Category Method Dataset Aggregation algorithm QHD = 0 QHD = 0.10 QHD = 0.17
Quantity distribution skew Min-size Dirichlet CIFAR10 FedAvg 5 3 2
Rand 5 3 1
FedProx 5 3 2
POC 5 3 2
MOON 6 3 2
Covtype FedAvg 3 2 1
Rand 3 2 1
FedProx 3 2 1
POC 3 2 1
MOON 4 2 1
Looking at Table VIII, it is clear that regardless of the degree of non-IIDness in the number of records from each label across clients, all aggregation algorithms converge after a consistent number of communication rounds. This phenomenon occurs due to the similarity in data distributions among clients, where each client’s dataset mirrors others, making any information obtained from all clients also obtainable from just one client.
V-D Spatiotemporal Skew
This section examines how varying levels of data disparity among clients, based on time and location, impact model performance.
Table IX: Mean and standard deviation Accuracy for each dataset for CL, FedAvg, Rand, FedProx, Power-Of-Choice, and MOON, considering different levels of non-IID partitioning of the records based on their space (SP) and time (TMP) measured by STHD for $K=30$. Each model has undergone five trials.
Category Type Method Dataset CL STHD FedAvg Rand FedProx POC MOON
SPT distribution skew SP skew St-Dirichlet Serengeti 96.95% $\pm$ 0.11% 0 95.58% $\pm$ 0.08% 95.56% $\pm$ 0.09% 95.62% $\pm$ 0.14 95.48% $\pm$ 0.09 96.69% $\pm$ 0.06
0.25 95.63% $\pm$ 0.05% 95.50% $\pm$ 0.10% 95.64% $\pm$ 0.03 95.51% $\pm$ 0.11 96.76% $\pm$ 0.08
0.5 95.40% $\pm$ 0.08% 95.32% $\pm$ 0.07% 95.36% $\pm$ 0.06 95.32% $\pm$ 0.06 96.51% $\pm$ 0.04
0.75 94.09% $\pm$ 0.25% 93.91% $\pm$ 0.36% 94.15% $\pm$ 0.21 94.06% $\pm$ 0.09 95.80% $\pm$ 0.16
0.9 88.65% $\pm$ 0.18% 87.68% $\pm$ 0.58% 88.49% $\pm$ 0.13 88.20% $\pm$ 0.39 91.74% $\pm$ 0.20
TMP skew 5GNTF 92.39% $\pm$ 0.10% 0 92.15% $\pm$ 0.04% 92.15% $\pm$ 0.02% 92.14% $\pm$ 0.02% 92.15% $\pm$ 0.04% 92.23% $\pm$ 0.02%
0.25 92.07% $\pm$ 0.12% 92.05% $\pm$ 0.15% 92.07% $\pm$ 0.18% 92.15% $\pm$ 0.05% 92.28% $\pm$ 0.04%
0.5 88.28% $\pm$ 1.87% 87.92% $\pm$ 2.00% 89.19% $\pm$ 2.09% 92.08% $\pm$ 0.08% 89.24% $\pm$ 1.80%
0.75 85.82% $\pm$ 0.06% 85.64% $\pm$ 0.08% 85.83% $\pm$ 0.08% 91.77% $\pm$ 0.32% 86.51% $\pm$ 1.79%
0.9 84.62% $\pm$ 0.36% 84.50% $\pm$ 0.20% 84.55% $\pm$ 0.37% 89.23% $\pm$ 2.13% 83.94% $\pm$ 0.02%
Number of times that performed the best 0 0 0 3 7
Table X: Number of rounds required to achieve 90% of the maximum accuracy reached accuracy levels of spatiotemporal non-IID cases as determined by STHD for K = 30
Dataset Aggregation algorithm STHD = 0 STHD = 0.25 STHD = 0.50 STHD = 0.75 STHD = 0.9
Serengeti FedAvg 6 7 7 10 14
Rand 7 7 8 10 14
FedProx 7 7 7 10 14
POC 7 7 7 10 15
MOON 8 8 9 12 19
V-D1 Classification power
In this subsection, we focus on the simulation results to evaluate various aggregation methods and datasets regarding classification accuracy. We pay particular attention to the impact of different levels of spatiotemporal skewness among the clients’ data.
Highlight 13: The higher the non-IIDness level in time and space, the worse the model’s performance.
Table IX demonstrates that irrespective of the aggregation algorithm employed, model performance deteriorates when data distribution among clients varies concerning time or space. This phenomenon occurs because increasing the differences among clients’ data based on time and location also raises the level of non-IIDness in the clients’ label distributions. This behavior is evident in Table XI, which displays the HD in label distributions at varying levels of STHD among clients. We also concluded in the Label Skew study section that higher levels of non-IIDness among clients’ data distributions negatively impact the performance of the final model.
V-D2 Convergence
In this subsection, we concentrate on our results related to the models’ convergence when there are variations in the data concerning time and location, considering the results obtained on Serengeti and 5GNTF datasets.
Highlight 14: The higher the non-IIDness level in time and space, the more required rounds to achieve convergence.
As shown in Table X, irrespective of the aggregation algorithm used for training, an increase in the distance between data distributions in terms of time and location results in a need for more rounds to achieve convergence. This trend is evident in both the Serengeti and 5GNTF datasets. When comparing the required rounds for STHD=0 and STHD=90, it is clear that in the worst case (STHD=90), the number of rounds needed is at least twice as high. This observation aligns with our earlier findings on classification performance, specifically from the Label Skew study, where we determined that higher non-IIDness among clients’ label distributions leads to increased rounds required to achieve convergence.
Table XI: HD among clients’ label distributions at varying levels of non-IID partitioning by time and space
Dataset HD = 0 HD = 0.25 HD = 0.50 HD = 0.75 HD = 0.9
Serengeti 0.01 0.09 0.22 0.36 0.53
5GNTF 0.03 0.20 0.29 0.30 0.49
V-E General results
In this section, we provide highlights summarizing the overall results obtained from our experiments, combining the behavior shown before for label, feature, quantity, and spatiotemporal skewness.
Highlight 15: FedProx, POC, and MOON demonstrate superior performance compared to FedAvg and Rand in general FL scenarios
Table XII: The number of cases in which each specific algorithm achieved the best performance for each study
Study #Cases FedAvg Rand FedProx POC MOON
Label Skew 20 3 1 8 2 6
Feature Skew 36 4 2 7 7 16
Quantity Skew 12 1 0 3 6 2
Spatio Temporal Skew 10 0 0 0 3 7
Total best performance 8 3 18 18 31
Table XII summarizes the cases in which each specific algorithm exhibited the best performance compared to other aggregation algorithms across four skewness types considered in our study. In most cases, the FedProx, POC, and MOON aggregation algorithms achieved the best performance, outperforming the simpler FedAvg and Rand algorithms.
Highlight 16: FedProx is more effective on image datasets, while MOON performs better with tabular datasets.
Table XIII: The number of cases in which each aggregation algorithm achieved the best performance for each type of dataset
Dataset type #Cases FedAvg Rand FedProx POC MOON
Image 34 7 3 15 9 0
Tabular 44 1 0 3 9 31
Total best performance 8 3 18 18 31
Table XIII examines the best-performing aggregation algorithms from the perspective of the dataset type used for training. It shows that FedProx outperforms all other algorithms on image datasets in fifteen out of thirty-four cases. In comparison, MOON generally surpasses other algorithms on tabular datasets in thirty out of forty-four cases.
VI Design Insights and Opportunities
We provide some design insights and opportunities intending to help researchers direct their efforts toward solving the effects of data heterogeneity.
Quantifying non-IIDness.
Several works claim that the data non-IIDness affects the performance of FL models [27, 29, 16]. Nevertheless, for the first time, we demonstrate that the effect of the non-IIDness is not the same under all the levels of heterogeneity (see Fig. 4). Therefore, it is vital to quantify the non-IIDness level in FL. This work uses the HD metric to measure the data heterogeneity. However, we encourage researchers to test different metrics such as Jensen-Shannon distance [33], Earth mover’s distance [9], and Total Variation distance [3], among others.
Better methods to tackle high non-IIDness,
This work showcases how the state-of-the-art methods to tackle non-IIDness (Rand, POC, FedProx, MOON) perform against FedAvg. The conclusion is that no algorithm works better than FedAvg in all the scenarios. Moreover, the methods that work better do not show a vast improvement regarding FedAvg under high non-IID scenarios as their gain is at most two percent points. This phenomenon has also been studied and claimed in the scarce work of empirical analysis of non-IID data and methodologies [43, 1, 31, 22]. Therefore, creating methods to appropriately alleviate the effect of high data heterogeneity is needed to evolve and preserve FL. The latter aligns with the open problems reported by Kairouz et al. [18].
Focusing on highly unbalanced data,
In our experiments, we claim that the more significant decrease in performance comparing CL and FL occurs in unbalanced datasets since it relates to the challenge of learning from highly skewed and less representative data (see Table II). Thus, it is relevant to create solutions to tackle non-IIDness by considering the degree of unbalancedness that the labels might have across the clients.
Studying spatiotemporal skew.
To the time of writing of this work, no analyses or empirical studies about the effect of the spatiotemporal skew on the performance of FL models exist. Thus, for the first time, we produce experiments to understand how different levels of spatiotemporal non-IIDness affect the prediction power of an FL model. The results show (see Table IX) that high levels of space or time skewness decreases the performance of the models, more specifically when the HD is higher than 0.75 (severe non-IIDness). Thus, researchers may benchmark techniques to deal with space and time skew in FL [40, 11, 49] to determine the behavior under high data heterogeneity levels.
Methods to compare mixed non-IIDness types.
Current tools and methods for synthetic partitioning centralized data into federated data [17, 47, 20, 34, 15] focus on simulating one type of non-IIDness (label, feature, quantity, spatiotemporal skewness). Nevertheless, a more realistic scenario would be combining two or more types of data heterogeneity to evaluate how those mixes can alter the performance of FL models. Therefore, for research in FL purposes, it would be interesting to create methods to partition centralized data into federated clients that permit the control of non-IIDness level for each non-IIDness type at the same time.
VII Conclusions
This study provides a comprehensive empirical analysis of the non-IID effect in FL. We benchmarked five state-of-the-art strategies for handling non-IID data distributions under controlled conditions, exploring label, feature, quantity, and spatiotemporal skewness, with a novel focus on the latter. We aim to standardize the methodology for studying data heterogeneity in FL by using HD to quantify data distribution differences. Our findings reveal the significant impact of label and spatiotemporal skew non-IID types on FL model performance. We also demonstrate that the model’s performance drop appears at a double threshold. When HD is higher than 0.5 and 0.75, a higher damage and a steeper decrease in performance slope occurs. Moreover, our results suggest that the FL performance is heavily affected mainly when the non-IIDness is extreme. Thus, we offer valuable recommendations for researchers to address data heterogeneity. This work represents the most thorough examination of non-IIDness in FL to date, providing a robust foundation for future research in FL.
VIII Acknowledgments
Daniel Mauricio Jimenez G. and Andrea Vitaletti were partially supported by PNRR351 TECHNOPOLE – NEXT GEN EU Roma Technopole – Digital Transition, FP2 – Energy transition and digital transition in urban regeneration and construction and Sapienza Ateneo Research grant “La disintermediazione della Pubblica Amministrazione: il ruolo della tecnologia blockchain e le sue implicazioni nei processi e nei ruoli della PA.” Aris Anagnostopoulos was supported by the ERC Advanced Grant 788893 AMDROMA, the EC H2020RIA project “SoBigData++” (871042), the PNRR MUR project PE0000013-FAIR, the PNRR MUR project IR0000013-SoBigData.it, and the MUR PRIN project 2022EKNE5K “Learning in Markets and Society.” Ioannis Chatzigiannakis was supported by PE07-SERICS (Security and Rights in the Cyberspace) – European Union Next-Generation-EU-PE0000014 (Piano Nazionale di Ripresa e Resilienza – PNRR). Andrea Vitaletti was supported by PE11 - MICS (Made in Italy – Circular and Sustainable) – European Union Next-Generation-EU (Piano Nazionale di Ripresa e Resilienza – PNRR) and the project SERICS (PE00000014) under the MUR National Recovery and Resilience Plan funded by the European Union - NextGenerationEU.